OpenStudy (anonymous):

Graph 2 periods of the equation f(x) = 3cot(x + pi/2)-1

OpenStudy (anonymous):

@thu1935 do you have any idea how to do this because I don't?

OpenStudy (anonymous):

To graph trig trig equation, set up a Sign Table (Sign Chart) with x varies from 0 to 2Pi (2 periods = 2Pi, since the period of cot is pi). Call f(x) = 3 cot(x + pi/2) - 1 Cal g(x) = cot (x + Pi/2). Set up a Sign Table (sign chart) Give x consecutive values from 0 to 2Pi.. Compute g(x) by a calculator. Compute related values of f(x). Remember: When x = Pi/2 and x = 3Pi/2, g(x) and f(x) = - infinity and + infinity (symbol: ||). The graph would be alike to the one in the graph, but with more computed details. Note: f(x) = 0 when cot(x + Pi/2) = 1/3 = cot A. Find A by calculator. Plot f(x) = 0 in the chart. x |0 Pi/4 Pi/2 3Pi/4 Pi 5Pi/4 3Pi/2 7Pi/4 2Pi --------------------------------------------------------- g(x)|0 - - || - - 0 - - II + + 0 --------------------------------------------------------- f(x)|-1 || -1 || -1

OpenStudy (anonymous):